Police officers confiscate a packet of white powder that they believe contains heroin. Purification by a forensic chemist yields a  41.11 mg sample for airproof combustion analysis. The sample gives  103.5 mg of CO2 and  22.11 mg of H2O. A second sample is analyzed for nitrogen content, which is 3.8%.


1
Expert's answer
2021-09-12T00:43:40-0400

Molar mass of CO2 = 44.01

= 103.5/44.01 = 2.35moles

Molar mass of H2O = 18.01528

22.11/18.01528 = 1.23 moles

1.23 + 2.35 = 3.58 moles

44.1 × 0.038= 1.68

= 1.68/3.58 = 0.47